find the equation of a line parallel to 2x-5y=10 and passes through the point (-2, -3) IN STANDARD FORM
Answer by jim_thompson5910(35256)

Anything parallel to 2x-5y=10 is in the form 2x-5y=C where C is some number.

Since the line passes through (-2, -3), we know that x = -2 and y = -3

Plug these values in and simplify

2x-5y=C

2(-2)-5(-3)=C

-4+15=C

11=C

So C = 11, which means that the answer is 2x-5y = 11


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim